# Monte Carlo simulations of relativistic shock breakout from a stellar wind

### v1

@@ C.2.2 On the Discrepancy in the Far Upstream Region
This temperature rise also leads to a significant discrepancy in $x_{l}$ between the simulation and the analytical model in the far upstream region, as observed across all models ($\Gamma_{u}=2$, 6, and 10). In this region, the simulation exhibits a significant rise in $x_{l}$ (taking place beyond the upstream boundary of the range displayed in Figures 9 - 11), followed by a relatively flat profile. This profile reflects the fact that the photon population is dominated by photons produced through the thermal bremsstrahlung process. In our simulation, the far-upstream condition is set as follows: the photons are in thermal equilibrium (blackbody distribution) with a very low temperature ($kT_{u}\ll m_{e}c^{2}$), and the photon-to-baryon ratio of the incoming flow from the upstream boundary is set to $n_{\gamma,u}/n_{u}=10^{-2}$. [^18] Specifically, $n_{\gamma,u}=10^{-2}n_{u}=10^{13}~{}{\rm cm}^{-3}$, which leads to $kT_{u}=k\left(\frac{n_{\gamma,u}c^{3}h^{3}}{16\pi\zeta(3)k^{3}}\right)^{1/3}\approx 0.68~{}{\rm eV},$ where $\zeta(3)\approx 1.202$ is the Riemann zeta function. However, note that the essential features of the shock structure do not depend on the details of the upstream condition as long as it remains sufficiently cold, ensuring that photons produced immediately downstream dominate over those from $n_{\gamma,u}$. This rapid heating drives the flow out of thermal equilibrium, leading to the rapid production of photons through thermal bremsstrahlung. While this process remains modest in the deceleration region, it has a noticeable effect in the far upstream region, significantly increasing the photon-to-baryon ratio $n_{\gamma}/n$ from its initial value.
@@
(Chapline & Stevens, 1973; Weaver, 1976), where $n_{15}=n/10^{15}~{}{\rm cm}^{-3}$. [^19] Since the emissivity of thermal bremsstrahlung scales as $f_{\nu}\propto\nu^{0}$ at $h\nu\lesssim kT$, the total (frequency-integrated) photon generation rate diverges unless a low-energy cutoff is specified. An effective cutoff naturally arises in the system, as absorption below a certain frequency prevents the photon number from increasing. Note that this cutoff is automatically incorporated in the simulation, as free-free absorption is explicitly taken into account. Note also that, the cutoff due to Coulomb screening, given in Equation (15), is irrelevant in this region, as its value is significantly lower due to the lower temperature and density. To a good approximation, $\Lambda_{ff}$ can be expressed as $\Lambda_{ff}\approx\max\left[1,\frac{1}{2}\ln(y)\{1.6+\ln(y)\}\right],$ where $y=kT/h\nu_{c}$ (Levinson & Nakar, 2020). In the range $\hat{T}=10^{-4}$ to $2\times 10^{-2}$, $\Lambda_{ff}$ varies from $\sim 1$ to $\sim 74$ for $n_{15}=1$.

### v2

@@ C.2.2 On the Discrepancy in the Far Upstream Region
This temperature rise also leads to a significant discrepancy in $x_{l}$ between the simulation and the analytical model in the far upstream region, as observed across all models ($\Gamma_{u}=2$, 6, and 10). In this region, the simulation exhibits a significant rise in $x_{l}$ (taking place beyond the upstream boundary of the range displayed in Figures 10 - 13), followed by a relatively flat profile. This profile reflects the fact that the photon population is dominated by photons produced through the thermal bremsstrahlung process. In our simulation, the far-upstream condition is set as follows: the photons are in thermal equilibrium (blackbody distribution) with a very low temperature ($kT_{u}\ll m_{e}c^{2}$), and the photon-to-baryon ratio of the incoming flow from the upstream boundary is set to $n_{\gamma,u}/n_{u}=10^{-2}$. [^18] Specifically, $n_{\gamma,u}=10^{-2}n_{u}=10^{13}~{\rm cm}^{-3}$, which leads to $kT_{u}=k\left(\frac{n_{\gamma,u}c^{3}h^{3}}{16\pi\zeta(3)k^{3}}\right)^{1/3}\approx 0.68~{\rm eV},$ where $\zeta(3)\approx 1.202$ is the Riemann zeta function. However, note that the essential features of the shock structure do not depend on the details of the upstream condition as long as it remains sufficiently cold, ensuring that photons produced immediately downstream dominate over those from $n_{\gamma,u}$. This rapid heating drives the flow out of thermal equilibrium, leading to the rapid production of photons through thermal bremsstrahlung. While this process remains modest in the deceleration region, it has a noticeable effect in the far upstream region, significantly increasing the photon-to-baryon ratio $n_{\gamma}/n$ from its initial value.
@@
(Chapline and Stevens, 1973; Weaver, 1976), where $n_{15}=n/10^{15}~{\rm cm}^{-3}$. [^19] Since the emissivity of thermal bremsstrahlung scales as $f_{\nu}\propto\nu^{0}$ at $h\nu\lesssim kT$, the total (frequency-integrated) photon generation rate diverges unless a low-energy cutoff is specified. An effective cutoff naturally arises in the system, as absorption below a certain frequency prevents the photon number from increasing. Note that this cutoff is automatically incorporated in the simulation, as free-free absorption is explicitly taken into account. Note also that, the cutoff due to Coulomb screening, given in Equation (15), is irrelevant in this region, as its value is significantly lower due to the lower temperature and density. To a good approximation, $\Lambda_{ff}$ can be expressed as $\Lambda_{ff}\approx\max\left[1,\frac{1}{2}\ln(y)\{1.6+\ln(y)\}\right],$ where $y=kT/h\nu_{c}$ (Levinson and Nakar, 2020). In the range $\hat{T}=10^{-4}$ to $2\times 10^{-2}$, $\Lambda_{ff}$ varies from $\sim 1$ to $\sim 74$ for $n_{15}=1$.